# Modeling the precipitation distribution by radius and pore size during drying of an impregnated sphere

### v1

@@ Abstract
The process of preparing heterogeneous catalysts on porous supports includes a drying stage, in which the porous material, impregnated with an aqueous solution of the catalyst precursor, is dried, and the precursor is precipitated on the pore walls. The precipitate distribution throughout the support volume strongly influences the catalyst performance and durability within industrial reactors. This paper presents a mathematical model for simulating the precipitation distribution during the drying of a porous sphere. Examples of calculations are given demonstrating the influence of some parameters on the distribution by space and by pore size.
@@ 1 Introduction
Heterogeneous catalysts are widely used in the modern chemical industry to accelerate the rate of reactions and change the selectivity of processes in the desired direction. To increase the efficiency of the catalyst, it is advisable to increase the contact area of its active elements with the reagents. This can be done by distributing the active elements over the pore surface in a porous support with a large internal surface area. One of the most common methods for preparing such catalysts is the impregnation method, in which the support is first impregnated with a precursor solution, followed by drying and subsequent calcination [1, 2]. In industrial reactors, such catalysts are used in the form of millimeter-sized pellets of various shapes, of which the sphere is the most convenient for modeling.
In the models for preparing supported catalysts by impregnation, two mechanisms for placing active elements on pore walls were considered. Chronologically, the first mechanism to be considered was one in which active elements were adsorbed from solution onto the pore walls. In this case, the distribution of active elements in the pore space was mainly formed at the impregnation stage, and subsequent drying only corrected this distribution [3, 4, 5, 6]. This mechanism works in cases where the concentration of the solution is low and the chemical interaction between the solute and solid support is strong enough. In another mechanism, the catalyst precursor in the form of a solt of catalytically active metal is deposited from the supersaturated aqueous solution on the pore walls during the drying step [7, 8].
In this paper, a mathematical model of the second mechanism of placement of the catalyst precursor in the pore space of the support is considered. The model under consideration, like model [7, 8], is based on the Whitaker model of drying porous materials [9, 10] with additional terms describing the precipitation of salt from solution. However, the equation for the gas phase is excluded from the equations of Model [9], since precipitation occurs only in the presence of a liquid solution in the pores. In this case, dynamic equilibrium between the liquid and vapor is usually assumed, i.e., the vapor density is equal to the saturated density. Under these conditions, the equation for the gas phase can be replaced by the requirement that the gas pressure in the pores be equal to the external gas pressure. This substitution has virtually no effect on the overall solution of the model.
The results of calculating the distribution of precipitant along the radius of a sphere in dependence on some model parameters are presented. To simulate precipitant distribution over pores, a capillary bundle model is used, which is used in the Whitaker model to estimate capillary pressure and fluid velocity within pores. It is assumed that sediment settles in those pores (capillaries) that are currently drying.
@@ 2.1 Equations
The mathematical description of the drying process of a solution in the pores of a material with precipitation presented here does not take into account mass and heat transfer in the air within the pore space. This is because, according to our computational experience, processes in the gas phase have virtually no effect on the distribution of precipitation. Precipitation occurs during the first stage of drying, when the liquid solution is still present in the pores of the material. The Whitaker drying model typically assumes a dynamic equilibrium between liquid and vapor in the pore space, but the presented model only works in the presence of liquid. Therefore, the equation for the gas phase is omitted, and it is assumed that the water vapor pressure in the pores is equal to the saturated vapor pressure at a given temperature.
The system of equations below describes the evolution of the fraction of the sphere’s volume occupied by the solution, $\epsilon_{l}(r,t)$, the mass fraction of the metal salt in the solution, $\gamma_{m}(r,t)$ ($\gamma_{w}=1-\gamma_{m}$), the fraction of the sphere’s volume occupied by the precipitate, $\epsilon_{m}(r,t)$, and the temperature $T(r,t)$ in the radial coordinate $r$ at time $t$. It is assumed that all phases are in thermal equilibrium and the temperature is the same in both the solid and liquid phases.
@@
Equations (1), (2) and (4) are continuity equations for water, dissolved salt, and heat (enthalpy), respectively, taking into account binary diffusion in solution and thermal conductivity. Equation (3) describes the accumulation of precipitated metal salt within the volume of the sphere. The precipitation rate, $r_{pr}$, is defined as
@@
where $k_{pr}$ is the rate constant, $\gamma_{m}^{*}$ is the mass fraction of salt in a saturated solution.
@@ 2.2 Parameters
so (assuming the water density is equal 1000 kg/m3) the limited nass fraction is defined as
@@
The surface tension and viscosity of a solution should depend on the salt concentration. However, information on these relationships is very scarce. Therefore, for the sake of certainty, it is assumed that in a saturated solution
@@
2.3 Velosity of liquid
The velocity of the liquid $\bm{v}_{l}$ is determined by Darcy’s law
@@ 2.3 Velosity of liquid
where $k_{l}$ is the permeability of porous material for liquid and $\mu_{l}$ is the dynamic viscosity of liqquid. The liquid pressure $P_{l}$ is definrd as
@@
where $P_{g}$ is the gas pressure in the pores, and $P_{c}$ is the capillary pressure of the liquid. Since the gas phase is not taken into account in the model, the gas pressure $P_{g}$ is assumed to be constant and equal to the gas pressure $P_{ex}$ in the drying environment.Therefore, in fact,
@@
The permeability of a porous material is determined by the structure of its pores. In this case, the density function $dV_{p}/dr_{p}$ of the pore volume $V_{p}$ distribution by their size (radius) $r_{p}$ plays an important role. Like many papers on this topic, we assume that the distribution density has the form of a truncated Gaussian density
@@
To calculate the velocity of a liquid, we use the model of a porous medium as a bundle of capillaries with volume distribution by radius (17). This model assumes that when the medium is soaked, the liquid fills the capillaries in order of increasing radius, and when the liquid dries, it fills them in the reverse order. Thus, the volume occupied by the liquid can be represented as follows:
@@
where $r_{p,f}$ is the largest radius of liquid-filled capillaries. For known $\epsilon_{l}$, formula (18) is considered as an equation for unknown $r_{p,f}$. By solving this equation, we find $r_{p,f}(\epsilon_{l})$ and calculate the pressure in a capillary with a moving liquid using Young-Laplace formula.
@@ 2.4 Boundary and initial conditions
The boundary conditions for system (1)-(4) are specified at the center of the sphere at $r=0$ and at the surface at $r=R$. At $r=0$, zero fluxes of the system variables are specified. At $r=R$, the flow of water is specified in the form of water vapor evaporating from the surface of the sphere:
@@
Initial conditions: the pores are completely filled with a solution with a uniform distribution of the mass fraction of salt and with a uniform distribution of temperature:
@@
At $t=0$ there is no sediment in the pores:
@@
is a stable stationary state of system (1)-(4) with the specified boundary conditions (21), (22). Due to the uniqueness theorem, the solution with the initial condition (23) will asymptotically approach the stationary state (25) at $t\to\infty$. (Strictly speaking, the solution will never dry completely.) Therefore, we stop solving the problem when the following condition is satisfied
@@ 2.5 Numerical method
Since most of the terms in equations (1)-(4) have a divergent form, it is natural to use the cuntrol volume method for the numerical solution of the system.
The sphere is divided into $N$ spherical layers by spherical surfaces of radius $r_{k}=(k/N)^{1/3}R$, $k=1,2,..,N$ (the results of calculations with $N$ = 100 are presented below). The volume of each layer equals to $V/N$, where $V$ is the volume of the spherical pellet. The equations (1-4) are integrated over each spherical layer, with the divergent terms being integrated exactly and the mean value theorem being used to estimate the integrals of the non-divergent terms. As a result of integration, a 4NN ODE system is obtained, which, after taking into account the boundary conditions, is solved using the MATLAB function ‘ode15s’.
@@ 3 Results
3.1 General view
Figure 1 presents a general outline of a particular solution of system (1)-(4). It is assumed that the sphere was impregnated with a salt solution with the mass fraction of salt = $\gamma_{m,0}$ = 0.2 at the temperature $T_{0}$ = 20°C in air at atmospheric pressure $P_{g,ex}$ = 101325 Pa with humidity $\varphi$ = 10%. Then the sphere and the surrounding air are quickly heated to a temperature of $T_{ex}$ = 40°C and maintained at this temperature until the solution dries, while the gas pressure and the density of water vapor do not change.
@@ 3.1 General view
In subplot (a), the space-time graph of $\epsilon_{l}(r,t)$ is shown. At the initial stage of drying (fast stage), for approximately 20 minutes, the volume of the solution in the pores quickly decreases due to the evaporation of water. Then comes the slow drying stage, which lasts for approximately 30 min, until condition (1) with $\delta$ = 0.001 is met. During the slow stage, the reduction in liquid volume occurs mainly due to the precipitation of salt.
The spatiotemporal plot of $\gamma_{m}(r,t)$ shown in subplot (b) shows that the salt mass fraction increases rapidly during the fast drying stage and then stabilizes and remains near the saturation value. Salt precipitation is shown in graph (c). It is evident that significantly more precipitation falls in the subsurface layers of the sphere than in the center. The reasons for this phenomenon are discussed in section 3.2. Finally, the temperature in graph (d) remains virtually constant throughout the drying process.
@@
The change in the mass of the substances involved in the process during drying is shown in Figure 2.
@@ 3.2 Details of precipitation
Figure 3 shows some details of the solution of system (1-4), explaining the features of the process of salt precipitation. Shown here are the graphs relating to the time $t_{*}$ = 15 min from the solution, the general form of which is shown in Fig. 1.
Figure 3: Some details of the solution of system (1-4) at time $t_{*}$ = 15 min.
From the graph $\epsilon_{l}(t_{*},r)$, shown in subplot (a), it follows that the volume occupied by the liquid decreases from the center of the sphere to its periphery. Therefore, the radius $r_{p,f}$ (eq. (18)) of the largest capillary filled with liquid also decreases and, accordingly, the capillary pressure $P_{c}$ (eq. (19)) increases. The graph of capillary pressure $P_{c}(t_{*},r)$ is shown in subplot (b) in relative units $P_{c}/P_{ex}$, where $P_{ex}$ is atmospheric pressure.
An increase in the function $P_{c}(r)$ leads to a positive liquid velocity $v_{l}(r)$, the graph of which is shown in subplot (c). The llquid flows from the center to the surface of the sphere, where water evaporates from the solution. The water evaporates, but the salt remains in solution, increasing the salt concentration at the surface of the sphere (subplot (d)). Therefore, the rate of precipitation is greater at the surface of the sphere than at the center.
3.3 Distribution of precipitation by radius
@@
Figure 4 shows the graphs of the distribution of the salt volume by the radius of the sphere after drying at a temperature of 40°C. The graphs are constructed based on the results of solving the system (1-4) with different initial concentrations of salt in the solution. with increasing mass of salt in the solution, the lines in the figure rise, maintaining their shape.
Figure 4: Distribution of precipitations by radius at the end of drying depending on the initial mass fraction of salt.
@@
Figure 5a shows that the final salt distribution is virtually independent of drying temperature. Despite the significant differences in drying time: 165 minutes at $T_{ex}$ = 20°C, 51 minutes at 40°C, and 20 minutes at 60°C, the salt distribution curves are almost identical. Note, however, that at all three temperatures $T_{ex}$ the value of the constant $k_{pr}$ was the same and equal to 50.
@@
Figure 5b shows the significant influence of the parameter $k_{pr}$ on the distribution of precipitated salt after the solution dries. As this constant increases, an increasing volume of salt precipitates in the near-surface region of the sphere and an increasingly smaller volume in its central region. If we assume that the parameter $k_{pr}$ increases with increasing temperature, then a similar dependence of the distribution on the drying temperature should be observed. As the temperature increases, more and more salt will precipitate near the surface of the sphere.

### v2

@@ Abstract
The process of preparing heterogeneous catalysts on porous supports includes a drying stage, in which the porous material, impregnated with an aqueous solution of the catalyst precursor, is dried, and the precursor is precipitated on the pore walls. The precipitate distribution throughout the support volume strongly influences the catalyst performance and durability within industrial reactors. This paper presents a mathematical model for simulating the precipitation distribution during the drying of a porous sphere. A method for calculating the precipitant distribution by pore size is proposed. Examples of calculations are given demonstrating the influence of some parameters on the distribution by space and by pore size.
@@ 1 Introduction
Heterogeneous catalysts are widely used in the modern chemical industry to accelerate reactions and alter process selectivity in the desired direction. Improving catalyst efficiency can be achieved, in particular, by increasing the contact area of its active elements with the reactants. This is achieved by distributing the active elements over the pore surface of a porous support with a large internal surface area. One of the most common methods for preparing such catalysts is impregnation, in which the support is first impregnated with a precursor solution, then dried and calcined [1, 2]. In industrial reactors, such catalysts are used in the form of millimeter-sized granules of various shapes, of which spherical ones are the most convenient for modeling.
In known models for preparing supported catalysts by impregnation, two mechanisms for depositing active elements on pore walls were considered. Chronologically, the first mechanism to be considered was one in which active elements were adsorbed from liquid onto the pore walls. In this case, the distribution of active elements in the pore space was mainly formed at the impregnation stage, and subsequent operations only corrected this distribution [3, 4, 5, 6]. This mechanism works in cases where the salt concentration in the liqud is low and the chemical interaction between the solute and solid support is strong enough. In another mechanism, the catalyst precursor in the form of a catalyst metal salt is deposited from the supersaturated aqueous solution on the pore walls during the drying step [7, 8].
In this paper, a mathematical model of the second mechanism of placement of the catalyst precursor in the pore space of the support is presented. This model, like model [7, 8], is based on the Whitaker model of drying porous materials [9, 10] with additional terms describing the precipitation of salt from liquid in the pore space. However, the equation for the gas phase is excluded from the equations of model [9], since precipitation occurs only in the presence of a liquid in the pores. In this case, dynamic equilibrium between the liquid and vapor is usually assumed, that is, the vapor density is equal to the saturated density. Under these conditions, the equation for the gas phase can be replaced by the requirement that the gas pressure in the pores be equal to the external gas pressure. This substitution has virtually no effect on the overall solution of the model.
The results of calculating the distribution of precipitant along the radius of a sphere in dependence on some model parameters are presented. To simulate precipitant distribution over pores, a capillary bundle model is used, which is used in the Whitaker model to estimate capillary pressure and fluid velocity within pores. It is assumed that precipitation settles in those pores (capillaries) that are currently drying.
@@ 2 Model
The mathematical model presented below for drying a liquid in a porous sphere is based on the Whitaker model [9] with additional terms that take into account the precipitation of the dissolved substance from a supersaturated solution.
@@ 2.1 Equations
The Whitaker model describes both the drying of liquid mobile moisture in pores and the drying of immobile moisture bound to the solid skeleton of the material. In this study, only the drying of mobile moisture – the solution from which the salt precipitates – is of interest. Therefore, the Whitaker system of equations is considered only in the presence of liquid in the pores. In this case, the solvability of the system does not require invoking the sorption isotherm of the material. It is sufficient to assume dynamic equilibrium between the liquid and water vapor in the pores. That is, assume that the vapor density $\rho_{v}$ is always equal to the saturated vapor density $\rho_{v}^{*}$.
The continuity equation for a liquid in the Whitaker system for an aqueous solution can be split into two equations: one for the aqueous portion of the liquid and one for the dissolved salt. The continuity equation for air in the pores is unimportant in the presence of liquid moisture and can be omitted, assuming that the gas pressure in the pores $P_{g}$ is constant and equal to the external pressure $P_{g,ex}$. The heat transfer equation for a sphere (ignoring the gas phase in the pores) remains in the system for modeling non-isothermal drying.
As a result of the simplifications and assumptions made from the system of governing equations presented in [7, 8], the system (1)-(4) below is obtained. This system describes the change in the fraction of the sphere’s volume occupied by the liquid, $\epsilon_{l}(r,t)$, the mass fraction of the metal salt in the liquid, $\gamma_{m}(r,t)$ ($\gamma_{w}=1-\gamma_{m}$), the fraction of the sphere’s volume occupied by the precipitate, $\epsilon_{m}(r,t)$, and the temperature $T(r,t)$ along the radial coordinate $r$ at time $t$. It is assumed that all phases are in thermal equilibrium, and the temperature of both the solid and liquid phases is the same.
@@
Equations (1) and (2) are the continuity equations for the aqueous and salt fractions of the liquid, respectively, taking into account binary diffusion in the liquid phase. Equation (3) describes the accumulation of immobile precipitated salt in the pores. The precipitation rate, $r_{pr}$, is defined as
@@
where $k_{pr}$ is the rate constant, $\gamma_{m}^{*}$ is the mass fraction of salt in a saturated solution. Equation (4) is the heat transfer equation in a three-component medium: solid skeleton (s), liquid (l), and precipitated salt (m), taking into account the flow of the liquid and the thermal conductivity of the medium.
@@ 2.2 Parameters
so (assuming the water density is equal 1000 kg/m3) the limited mass fraction is defined as
@@
The surface tension and viscosity of a liquid should depend on the salt concentration. However, information on these relationships is very scarce. Therefore, for the sake of certainty, it is assumed that in a saturated solution
@@
2.3 Velocity of liquid
The velocity $\bm{v}_{l}$ of the liquid is determined by Darcy’s law
@@ 2.3 Velocity of liquid
where $k_{l}$ [m2] is the permeability of porous material for liquid and $\mu_{l}$ is the dynamic viscosity of liquid. The liquid pressure $P_{l}$ is definrd as
@@
where $P_{g}$ is the gas pressure in the pores, and $P_{c}$ is the capillary pressure of the liquid. Since the gas phase is not taken into account in the model, the gas pressure $P_{g}$ is assumed to be constant and equal to the gas pressure $P_{g,ex}$ in the drying environment.Therefore, in fact,
@@
The permeability of a porous material is determined by the structure of its pores. In this case, the density function $dV_{p}/dr_{p}$ of the pore volume $V_{p}$ distribution by pore size (radius) $r_{p}$ plays an important role. Like many papers on this topic, it is assumed that the distribution density has the form of a truncated Gaussian density
@@
To calculate the velocity of a liquid, the model of a porous medium as a bundle of capillaries is used with the volume distribution by radius (17). This model assumes that when the medium is soaked, the liquid fills the capillaries in order of increasing radius, and when the liquid dries, it leaves them in the reverse order. Thus, the volume occupied by the liquid can be represented as follows:
@@
where $r_{p,f}$ is the largest radius of liquid-filled capillaries. For known $\epsilon_{l}$, formula (18) is considered as an equation for unknown $r_{p,f}$. By solving this equation, one can find $r_{p,f}(\epsilon_{l})$ and calculate the pressure in a capillary with a moving liquid using Young-Laplace formula.
@@ 2.4 Boundary and initial conditions
The boundary conditions for equations (1), (2), (4) are specified at the center of the sphere at $r=0$ and at the surface at $r=R$. At $r=0$, zero fluxes of the system variables are specified. At $r=R$, the flow of water is specified in the form of water vapor evaporating from the surface of the sphere:
@@
where $\beta$ [m/s] is the external mass transfer coefficient ($\beta$ = 0.015 [7]), $M_{v}$ (= 18 g/mol) is the molar weight of water vape, $R_{g}$ is the ideal gas constant. The flow of $\gamma_{m}$ at $r=R$ is zero.
@@
$T_{g,b}(t)=\min(T_{0}+k_{b}t,T_{g,ex}),\;\mathrm{at}\,T_{0}\leq T_{g,ex},$ (23)
where $k_{b}$ is the rate of temperature change. In this paper, $k_{b}$ = 5 °C/s.
Initial conditions: the pores are completely filled with a liquid with a uniform distribution of the mass fraction of salt and with a uniform distribution of temperature:
@@
At $t=0$ there is no precipitant in the pores:
@@
is a stable stationary state of system (1)-(4) with the specified boundary conditions (21), (22). Due to the uniqueness theorem, the solution with the initial condition (24) will asymptotically approach the stationary state (26) at $t\to\infty$. (Strictly speaking, the liquid in the sphere will never dry completely.) Therefore, solving the problem is stopped when the following condition is satisfied
@@ 2.5 Numerical method
Since most of the terms in equations (1)-(4) have a divergent form, it is natural to use the control volume method for the numerical solution of the system.
To implement this method, the sphere is divided into $N$ spherical layers by spherical surfaces of radius $r_{k}=(k/N)^{1/3}R$, $k=1,2,..,N$ (the results of calculations with $N$ = 100 are presented below). The volume of each layer equals to $V/N$, where $V$ is the volume of the spherical pellet. The equations (1)-(4) are integrated over each spherical layer, with the divergent terms being integrated exactly and the mean value theorem being used to estimate the integrals of the non-divergent terms. As a result of integration, a 4NN ODE system is obtained, which, after taking into account the boundary conditions, is solved using the MATLAB function ‘ode15s’. All results presented below were obtained with a relative tolerance RelTol = 10-6 and an absolute tolerance AbsTol = 10-9.
@@ 3 Results
3.1 General view of the model solution
A general overview of a typical solution to system (1)–(4) is presented in Figure 1. This figure shows the space-time graphs of four system variables for solving the initial-boundary value problem under the following conditions. The sphere was impregnated with a liquid at a temperature of $T_{0}$ = 20°C. The mass fraction of salt in the liquid is $\gamma_{m,0}$ = 0.2. Drying occurs in air at atmospheric pressure ($P_{g,ex}$ = 101325 Pa) and a temperature of $T_{ex}$ = 40°C. The water vapor pressure in the drying air is constant and equal to $P_{v,ex}$ = 2333 Pa (which corresponds to 10% air humidity at a temperature of 20°C). This pressure was the same in all the examples below. The numerical solving of the problem was terminated when condition (27) with $\delta$ = 0.001 was satisfied. Let us consider the behavior of each variable during drying.
Subplot (a). A graph of $\epsilon_{l}(r,t)$ is shown. During the initial drying stage, over the course of approximately 20 minutes, the fraction of the sphere’s volume occupied by liquid rapidly decreases from an initial value of 0.6 to approximately 0.05 due to water evaporation and salt precipitation. This is followed by a slow drying stage, which lasts approximately 30 minutes, until condition (1) is satisfied. At each $t$, $\epsilon_{l}$ decreases monotonically with $r$. At the surface of the sphere, $\epsilon_{l}$ is slightly smaller than at the center.
@@ 3.1 General view of the model solution
Subplot (b). A graph of the mass fraction of salt $\gamma_{m}(r,t)$ in the liquid is shown. During the first 5 minutes of drying, the mass fraction of salt increases from an initial value of 0.2 to a saturation value of 0.29 at a given temperature $T_{ex}$ = 40°C and then remains virtually unchanged until the end of drying. Moreover, at the surface of the sphere $\gamma_{m}$ is slightly larger than in the center.
Subplot (c).A graph of the fraction of the sphere’s volume $\epsilon_{m}(r,t)$ occupied by salt precipitation is shown. It is clear that significantly more precipitation falls in the sphere’s surface layers than in its center. The reasons for this phenomenon are discussed in Section 3.2..
Subplot (d). The temperature graph $T(r,t)$ is shown. The temperature remains virtually constant throughout the drying process, with the exception of a jump from $T_{0}$ to $T_{ex}$ at the very beginning of the process.
The total change in the mass of the components involved in the process during drying is shown in Figure 2.
@@ 3.2 Details of precipitation
Figure 3 shows some details of the solution of system (1)-(4), explaining the features of the process of salt precipitation. Shown here are the graphs relating to the time $t_{*}$ = 15 min from the problem solution, the general form of which is shown in Fig. 1.
Figure 3: Details of the solution of system (1)-(4) at time $t_{*}$ = 15 min.
From the graph of $\epsilon_{l}(t_{*},r)$, shown in subplot (a), it follows that the volume occupied by the liquid decreases from the center of the sphere to its periphery. Therefore, the radius $r_{p,f}$ (eq. (18)) of the largest capillary filled with liquid also decreases and, accordingly, the capillary pressure $P_{c}$ (eq. (19)) increases. The graph of capillary pressure $P_{c}(t_{*},r)$ is shown in subplot (b) in relative dimensionless units $P_{c}/P_{ex}$, where $P_{ex}$ is the atmospheric pressure.
An increase in the function $P_{c}(r)$ leads to a positive liquid velocity $v_{l}(r)$, the graph of which is shown in subplot (c). The llquid flows from the center to the surface of the sphere, where water evaporates from the liquid. The water evaporates, but the salt remains in liquid, increasing the salt concentration at the surface of the sphere (subplot (d)). Therefore, the rate of precipitation is greater at the surface of the sphere than at the center.
3.3 Precipitation distribution over radius
@@
Figure 4 shows the graphs of the distribution of the salt volume by the radius of the sphere after drying at a temperature of 40°C. The graphs are constructed based on the results of solving the system (1)-(4) with different initial concentrations of salt in the aqua. With increasing mass of salt in the aqua, the lines in the figure rise, maintaining their shape.
Figure 4: Distribution of precipitation by radius at the end of drying depending on the initial mass fraction of salt.
@@
Figure 5a shows that the final salt distribution is virtually independent of drying temperature. Despite the significant differences in drying time: 165 minutes at $T_{ex}$ = 20°C, 52 minutes at 40°C, and 20 minutes at 60°C, the salt distribution curves are almost identical. Note, however, that at all three temperatures $T_{ex}$ the value of the constant $k_{pr}$ was the same and equal to 50.
@@
Figure 5b shows the significant influence of the parameter $k_{pr}$ on the distribution of precipitated salt after the liquid dries. As this constant increases, an increasing volume of salt precipitates in the near-surface region of the sphere and an increasingly smaller volume in its central region. If one assumes that the parameter $k_{pr}$ increases with increasing temperature, then a similar dependence of the distribution on the drying temperature should be observed. As the temperature increases, more and more salt will precipitate near the surface of the sphere.